OpenStudy (anonymous):

simplify the expression 6(2x+4)

jigglypuff314 (jigglypuff314):

distributive property is a(b+c) = ab + ac for example: 3(4x+2) = 3(4x) + 3(2) = 12x + 6 so 6(2x+4) = ?

OpenStudy (anonymous):

so it would be 12x+4

OpenStudy (anonymous):

12x+24

jigglypuff314 (jigglypuff314):

correct :)

OpenStudy (anonymous):

which one the 12x+4 or the 12x+24

jigglypuff314 (jigglypuff314):

12x+24 because 6(2x+4) = 6(2x) + 6(4) = 12x + 24

OpenStudy (anonymous):

ok thanks was confused if i should multiply both u really helped me with this problem

jigglypuff314 (jigglypuff314):

glad I could help :)

OpenStudy (anonymous):

i have some more problems i could use some help on if ur not busy

jigglypuff314 (jigglypuff314):

I'll be more than willing to help :)

OpenStudy (anonymous):

Select the product of (9x - 4)(4x + 8)

OpenStudy (anonymous):

i kno product means multiply

jigglypuff314 (jigglypuff314):

This would be the foil method FOIL = First Outer Inter Last for example: (3x-2)(2x+1) = (3x)(2x)+(3x)(1)+(-2)(2x)+(-2)(1) = 6x² + 3x - 4x - 2 = 6x² - x - 2

OpenStudy (anonymous):

oh ok ill have to keep that FOIL thing in mind

jigglypuff314 (jigglypuff314):

so (9x - 4)(4x + 8) becomes (9x)*(4x) + (9x)*(8) + (-4)*(4x) + (-4)(8) = ?